OpenStudy (shaeelynn):

What is the sum of the finite arithmetic series? (-5)+0+5+10+...+65 a)900 b)455 c)450 d)445

OpenStudy (aveline):

http://www.regentsprep.org/regents/math/algtrig/atp2/arithseq.htm Scroll down to find the equation to find the sum of the finite series.

ganeshie8 (ganeshie8):

65 = 13*5 so there are 13+2 = 15 terms in the series

ganeshie8 (ganeshie8):

first term is -5 last term is 65

ganeshie8 (ganeshie8):

use the series formula : sum = n/2(first term + last term)

ganeshie8 (ganeshie8):

here n = number of terms = 15

OpenStudy (shaeelynn):

i got 120 tho??

ganeshie8 (ganeshie8):

15/2(-5 + 65)

ganeshie8 (ganeshie8):

15/2(60) 15(30) 450

OpenStudy (shaeelynn):

thank youu